OpenStudy (anonymous):

Given f(x) = x^3+9x^2-5x-33 , find the sum of the squares of its zeros

OpenStudy (perl):

graph it, find the x intercepts, then you can take the sum of its squares

OpenStudy (anonymous):

what does it mean sum of squares

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (perl):

first find the x intercepts

OpenStudy (perl):

you can use wolfram

OpenStudy (paxpolaris):

if the zeroes are a, b and c find\(a^2+b^2+ c^2\)

OpenStudy (anonymous):

we were only allowed to use the sums and product

OpenStudy (anonymous):

sum is -9 and product is -33

OpenStudy (rsadhvika):

use Vieta's formulas and Newton sums

OpenStudy (paxpolaris):

\[(a+b+c)^2 \\ = \left[ \left( a+b \right) +c\right]^2\\ =\left( a^2+b^2+2ab \right)+c^2+2c(a+b)\\=a^2+b^2+c^2+2(ab+ac+bc)\]

OpenStudy (rsadhvika):

Nice :)

OpenStudy (paxpolaris):

\[a^2 +b^2+c^2=(a+b+c)^2-2(ab+ac+bc)\] .... not sure if this helps though?

OpenStudy (anonymous):

Yes it does @PaxPolaris thank you!

OpenStudy (paxpolaris):

how do you know (ab+ac+bc) @Shruti_ ?

OpenStudy (anonymous):

when you foil and factor

OpenStudy (anonymous):

what would you substitute in for a b and c

OpenStudy (paxpolaris):

a+b+c = -9

OpenStudy (anonymous):

right but what about the second half

OpenStudy (rsadhvika):

\[\large f(x) = c_3x^3 + c_2x^2 + c_1x^1 + c_0\] \(\large a+b+c = -\dfrac{c_2}{c_3}\) \(\large ab+bc+ca = \dfrac{c_1}{c_3}\) \(\large abc =- \dfrac{c_0}{c_3}\)

OpenStudy (paxpolaris):

how about -33/c + -33/b + -33/a

OpenStudy (anonymous):

thats what I got! where would i go from there?

OpenStudy (paxpolaris):

dunno :s

OpenStudy (anonymous):

Thanks for your help! @PaxPolaris

OpenStudy (rsadhvika):

\[a^2 +b^2+c^2=(a+b+c)^2-2(ab+ac+bc) = (-9)^2 -2(-5)^2\] simplify and you're done right ?

OpenStudy (anonymous):

How would it be -5? @rsadhvika

OpenStudy (rsadhvika):

VIeta's formula

OpenStudy (anonymous):

ohhhh okay thank you for elaborating!! @rsadhvika

OpenStudy (perl):

@rsadhvika how did you get that ?

OpenStudy (rsadhvika):

deriving them is easy, just compare the coefficients of the given polynomial and the factored form

OpenStudy (rsadhvika):

say a,b,c are the zeroes of given polynomial : \[\large x^3+9x^2-5x-33 = (x-a)(x-b)(x-c)\]

OpenStudy (rsadhvika):

compare the coefficients of x^2 term both sides you will get : \(\large 9 = -a-b-c\)

OpenStudy (rsadhvika):

compare the coefficients of x term both sides you get : \(\large -5 = ab+bc+ca\)

OpenStudy (rsadhvika):

similarly comparing constant terms you get \(\large -33 = -abc\)

OpenStudy (perl):

you can actually solve for a,b,c from those three equations

OpenStudy (rsadhvika):

Indeed.

OpenStudy (perl):

ok but someone above said to use a different formula

OpenStudy (paxpolaris):

but a,b and c look crazy ... the sum of squares is a nice integer @perl

OpenStudy (rsadhvika):

true, thats the reason we use newton sums

OpenStudy (rsadhvika):

newwton sums is very useful for finding sums of higher powers

OpenStudy (paxpolaris):

so 91 your answer @Shruti_

OpenStudy (anonymous):

Yes!